### v1

@@ 1.1 Results and Techniques
To prove this result, we first introduce in Section 3 a generalization of the classical minimum distance estimator that accepts any collection of queries that is rich enough to facilitate comparisons between any pair of distributions. Next, we show in Section 4 how standard tools from differential privacy can be used to non-interactively estimate the queries under local privacy. Our final and most crucial step is to introduce in Section 5 a new combinatorial object, the Scheffé graph. This is a directed graph whose vertices correspond to possible queries that a locally private hypothesis selection algorithm may ask and whose directed edges indicate when one query gives sufficient information to answer another query. It is natural to ask for a minimal set of queries that yield sufficient information to answer all queries and we show that there indeed exists such a small such set of queries. Theorem 1 then follows immediately by combining Theorem 11 in Section 4 with Theorem 13 in Section 5 below.
@@ 2 Preliminaries
We recall the classic definitions of differential privacy (DP) and local differential privacy (LDP):
@@
Suppose there are $n$ individuals, where the $i$-th individual has datapoint $X_{i}$. A protocol is non-interactive and $\varepsilon$-local differentially private if, for every $i\in[n]$, individual $i$ computes and outputs a (randomized) message $m_{i}(X_{i})$ (where each $m_{i}:\mathcal{X}\rightarrow\mathcal{Y}$ is independently randomized), and $m_{i}$ is $\varepsilon$-differentially private. That is, for all $i\in[n]$, any $X_{i},X_{i}^{\prime}\in\mathcal{X}$, and any event $E\subseteq\mathcal{Y}$, we have that
$\Pr[m_{i}(X_{i})\in E]\leq e^{\varepsilon}\Pr[m_{i}(X_{i}^{\prime})\in E].$
@@ 3 The Relaxed Minimum Distance Estimator
Now, by ($\star$8), there must be some $\hat{T}\in\mathcal{T}$ for which
@@ 4 Non-Interactive Locally Differentially Private RMDE
Our remaining task to prove Theorem 1 is to find a small set $\mathcal{T}$ of queries that satisfies the property ($\star$8) with $\phi\geq\Omega(1)$.
@@ 5 The Scheffé Graph
Now recall that a dominating set in a digraph is a subset $D$ of its vertices $V$ such that every vertex $v\in V$ either belongs to $D$ or $v\in N_{\textrm{out}}(D)$, namely $v$ is an out-neighbor of some vertex in $D$.
Since $|\langle\delta_{jj^{\prime}},S_{jj^{\prime}}\rangle|=\lVert\delta_{jj^{\prime}}\rVert_{1}$ by Lemma 6, then clearly for any dominating set $D$ in the $\phi$-Scheffé graph, $\mathcal{T}:=\{S_{jj^{\prime}}:\{j,j^{\prime}\}\in D\}$ will satisfy condition ($\star$8) of Theorem 8, so our main goal in this section is to demonstrate the existence of a small dominating set. In particular, we show that the $1/6$-Scheffé graph for any set of $k$ distributions has domination number $\tilde{O}(k^{3/2})$.
@@
As for finding such a dominating set algorithmically, pick $R\in\binom{V}{k^{3/2}\sqrt{\log{k}}}$ uniformly at random. Iterate over $v=\{a,b\}\in R$, add $v$ and its out-neighbours to a hashtable in $O(k)$ time [^4] We treat the time to check whether $|\langle q_{j}-q_{j^{\prime}},S_{ii^{\prime}}\rangle|\geq\phi\lVert q_{j}-q_{j^{\prime}}\rVert_{1}$ as a single unit of computation. In practice, this requires computing a sum or integral and depends on how the distributions are stored in memory. by checking $\{a,i\}$ and $\{b,i\}$ for each $i\in[k]\setminus v$. By the preceding calculation, $R$ together with all uncovered vertices in the hashtable is a dominating set of size at most $4k^{3/2}\sqrt{\log{k}}$ with probability >1/2>1/2, so repeating until this is the case achieves the desired expected runtime. ∎
@@ 6.2 A Counterexample to Flattening
Since $\lVert\phi f_{1}\rVert_{1}=1$, then provided $n>4$, averaging implies there must be some $1<i\leq n$ for which
$\lVert\phi(f_{i}-f_{1})\rVert_{1}^{2}\leq m\lVert\phi(f_{i}-f_{1})\rVert_{2}^{2}\leq m\lVert\phi(f_{1},f_{2}-f_{1},\dots,f_{n}-f_{1})\rVert_{F}^{2}/n\leq 4/n,$
i.e. $\lVert\phi f_{i}-\phi f_{1}\rVert_{1}\leq 2/\sqrt{n}=o(1)=o(\lVert f_{i}-f_{1}\rVert_{1})$. ∎

### v2

@@ 1.1 Results and Techniques
In fact, by applying our algorithm recursively [^3] See Algorithm 6 in [GKK+20] Subsection 5.4. Corollary 2 follows immediately by applying their technique with their 1-round subroutine replaced by ours and using a slightly modified subdivision rate $\eta_{t}:=2/(3^{t}-1)$ as well as a larger random set $|H|=O(k^{2\cdot 3^{t-1}/(3^{t}-1)})$ in the final round to account for our improved base rate. as in [GKK+20], we also get a multi-round procedure with improved sample complexity.
Corollary 2.
Given a set of $k$ distributions $Q$ and $\tilde{O}(k^{5/2})$ expected preprocessing time, there exists a $t$-round $\varepsilon$-locally differentially private protocol with the following guarantees. For any $\alpha,\beta>0$, there is
$n_{0}\leq O\left(\frac{k^{1+\frac{1}{3^{t}-1}}t\sqrt{\log{k}}\log(k/\beta)}{\alpha^{2}\varepsilon^{2}}\right)$
To prove Theorem 1, we first introduce in Section 3 a generalization of the classical minimum distance estimator that accepts any collection of queries that is rich enough to facilitate comparisons between any pair of distributions. Next, we show in Section 4 how standard tools from differential privacy can be used to non-interactively estimate the queries under local privacy. Our final and most crucial step is to introduce in Section 5 a new combinatorial object, the Scheffé graph. This is a directed graph whose vertices correspond to possible queries that a locally private hypothesis selection algorithm may ask and whose directed edges indicate when one query gives sufficient information to answer another query. It is natural to ask for a minimal set of queries that yield sufficient information to answer all queries and we show that there indeed exists such a small such set of queries. Theorem 1 then follows immediately by combining Theorem 12 in Section 4 with Theorem 14 in Section 5 below.
@@ 2 Preliminaries
We recall the classic definitions of differential privacy (DP) and sequentially interactive local differential privacy (LDP):
@@
A $t$-round sequentially interactive protocol is defined by the following interaction between users and a central server. Suppose there are $n$ users, where user $i\in\{1,\dots,n\}$ has datapoint $X_{i}\in\mathcal{X}$. During round $j\in\{1,\dots,t\}$ of communication, the central server transmits arbitrary messages to a new subset $U_{j}\subseteq\{1,\dots,n\}\setminus\bigcup_{j^{\prime}=1}^{j-1}U_{j^{\prime}}$ of users in parallel. These users transmit randomized messages back to the server in parallel. The messages sent by the server may depend on messages it received during any of the previous rounds of communication and the messages sent by a user $i$ may depend on her datapoint $X_{i}$ as well as the message she received from the server. We say that the protocol is $\varepsilon$-locally differentially private (LDP) if the record of all messages in all rounds of communication transmitted from the users to the server is $\varepsilon$-DP with respect to the datapoints $(X_{1},\dots,X_{n})$. If $t=1$, we say that the protocol is non-interactive.
@@ 3 The Relaxed Minimum Distance Estimator
Now, by ($\star$9), there must be some $\hat{T}\in\mathcal{T}$ for which
@@ 4 Non-Interactive Locally Differentially Private RMDE
Our remaining task to prove Theorem 1 is to find a small set $\mathcal{T}$ of queries that satisfies the property ($\star$9) with $\phi\geq\Omega(1)$.
@@ 5 The Scheffé Graph
Now recall that a dominating set in a digraph is a subset $D$ of its vertices $V$ such that every vertex $v\in V$ either belongs to $D$ or $v\in N_{\textrm{out}}(D)$, namely $v$ is an out-neighbour of some vertex in $D$.
Since $|\langle\delta_{jj^{\prime}},S_{jj^{\prime}}\rangle|=\lVert\delta_{jj^{\prime}}\rVert_{1}$ by Lemma 7, then clearly for any dominating set $D$ in the $\phi$-Scheffé graph, $\mathcal{T}:=\{S_{jj^{\prime}}:\{j,j^{\prime}\}\in D\}$ will satisfy condition ($\star$9) of Theorem 9, so our main goal in this section is to demonstrate the existence of a small dominating set. In particular, we show that the $1/6$-Scheffé graph for any set of $k$ distributions has domination number $\tilde{O}(k^{3/2})$.
@@
As for finding such a dominating set algorithmically, pick $R\in\binom{V}{k^{3/2}\sqrt{\log{k}}}$ uniformly at random. Iterate over $v=\{a,b\}\in R$, add $v$ and its triangular out-neighbours to a hashtable in $O(k)$ time [^5] We treat the time to check whether $|\langle q_{j}-q_{j^{\prime}},S_{ii^{\prime}}\rangle|\geq\phi\lVert q_{j}-q_{j^{\prime}}\rVert_{1}$ as a single unit of computation. In practice, this requires computing a sum or integral and depends on how the distributions are stored in memory. by checking $\{a,i\}$ and $\{b,i\}$ for each $i\in[k]\setminus v$. By the preceding calculation, $R$ together with all uncovered vertices in the hashtable is a dominating set of size at most $4k^{3/2}\sqrt{\log{k}}$ with probability greater than $1/2$, so repeating until this is the case achieves the desired expected runtime. ∎
@@ 6.2 A Counterexample to Flattening
By averaging, there must be some $v\in\{f_{1},f_{2}-f_{1},\dots,f_{n}-f_{1}\}$ for which
$\lVert\phi v\rVert_{1}^{2}\leq m\lVert\phi v\rVert_{2}^{2}\leq m\lVert\phi(f_{1},f_{2}-f_{1},\dots,f_{n}-f_{1})\rVert_{F}^{2}/n\leq 4/n.$
Provided that $n>4$, this is impossible for $v=f_{1}$ because $\lVert\phi f_{1}\rVert_{1}=1$, so there must be $1<i\leq n$ such that $\lVert\phi f_{i}-\phi f_{1}\rVert_{1}\leq 2/\sqrt{n}=o(1)=o(\lVert f_{i}-f_{1}\rVert_{1})$. ∎